3w3 − 24w2 + 45w = 0
w=
w=
w=

Answer by algebrahouse.com(1659)

3w³ - 24w² + 45w = 0
3w(w² - 8w + 15) = 0 {factored out the greatest common factor, 3w}
3w(w - 5)(w - 3) = 0 {factored parentheses into two binomials}
3w = 0 or w - 5 = 0 or w - 3 = 0 {set each factor equal to 0}
w = 0 or w = 5 or w = 3 {solved each equation for w}

For more help from me, visit: www.algebrahouse.com